OpenStudy (anonymous):

Find the derivative of the function y = 2x^2 - 13x + 5 and use it to find the equation of the line tangeant to the curve at x = 3

OpenStudy (anonymous):

y'=4x - 13 and at x = 3 y'(3) = -1

OpenStudy (anonymous):

Can you show me how you did it?

OpenStudy (anonymous):

the derivative?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

y'=n x^(n-1) right? and the derivative of any constant is 0 so the 5 at the end of the equation equals 0 so, y = 2*2 x^(2-1) - 13 x^(1-1) +0 = 4x - 13

OpenStudy (anonymous):

OHHHHHHHHHHH You just jump started my memory from last year's math class, thanks.

OpenStudy (anonymous):

haha np man

OpenStudy (anonymous):

do you reckon you can help with my question?

OpenStudy (anonymous):

Uhhh I might what is it?